The floatref package^{∗}

Andrew D. Lewis

1997/03/09

1 How to use the floatref macros

\figref This package is really trivial. It defines two macros, \figref and \tabref. They \tabref each do the same thing, \figref for figures, and \tabref for tables. To refer to

a figure labelled with label type \figref{label} and similarly for \tabref. The point is that “Figure” is automatically written so you don’t have to write it every time.

2 Implementation

1 ⟨∗package⟩

2

 3 \newcommand{\figref}[1]{\mbox{\ignorespaces\figurename~\ref{#1}\unskip}}
 4 \newcommand{\tabref}[1]{\mbox{\ignorespaces\tablename~\ref{#1}\unskip}}

5 ⟨/package⟩

∗This file has version number v1.0, last revised 1997/03/09.